xy"+2y'=0 reduced into y


1
Expert's answer
2021-07-04T17:27:25-0400

Let us solve "xy''+2y'=0." Let "z=y'." Then "xz'+2z=0." It follows that "x\frac{dz}{dx}=-2z," and hence "\frac{dz}{z}=-2\frac{dx}{x}." We have that "\int\frac{dz}{z}=-2\int\frac{dx}{x}," and thus "\ln|z|=-2\ln|x|+\ln|C|=\ln|\frac{C}{x^2}|" . We conclude that "z=\frac{C}{x^2}." Then "y'=\frac{C}{x^2}," and hence the solution is "y=-\frac{C}{x}+C_2."